OpenStudy (anonymous):

Solve the IVP: y''+y=0, y(pi/3)=0, y'(pi/3)=2. Find y(pi/6).

zepdrix (zepdrix):

So this is linear and homogeneous. Do you understand how to find the characteristic equation?

zepdrix (zepdrix):

Oh also, \(\Large \color{#3366CF}{\text{Welcome to OpenStudy! :)}}\)

OpenStudy (anonymous):

Thanks

zepdrix (zepdrix):

We assume our solution is of the form: \(\Large y=e^{rx}\) Taking it's derivative twice gives us: \(\Large y''=r^2e^{rx}\) Plugging these into our equation gives us:\[\Large r^2e^{rx}+e^{rx}=0\]Factoring gives us our characteristic equation:\[\Large e^{rx}(r^2+1)=0, \qquad\to\qquad r^2+1=0\]

zepdrix (zepdrix):

Hmm so it looks like we end up with complex roots for the characteristic equation. Do you remember what type of solutions we get when our roots are complex? :o

OpenStudy (anonymous):

No

zepdrix (zepdrix):

Roots of the form: \(\Large r\quad=\quad \alpha\pm\beta i\) Give us a solution of the form: \(\Large y\quad=\quad c_1 e^{\alpha x}\cos\beta x+c_2e^{\alpha x}\sin\beta x\)Exponentials AND sines/cosines. In our case, we have no real portion as part of our roots, so we won't get the exponentials.\[\Large r=\pm i\]Is what we got.

zepdrix (zepdrix):

So our beta is 1.\[\Large y(x)\quad=\quad c_1\cos x+c_2 \sin x\]To solve for our unknown constants, we need to use the initial conditions that were given.

zepdrix (zepdrix):

Oh I guess we need to find y'(x) at this point :o Do you understand how to find the derivative here? :D

zepdrix (zepdrix):

Too confusing? :c

OpenStudy (anonymous):

DEs are a struggle for me

zepdrix (zepdrix):

\[\Large y'(x)\quad=\quad ?\]So we ignore the constants. Derivative of cosine?

OpenStudy (anonymous):

-sin

zepdrix (zepdrix):

\[\Large y'(x)\quad=\quad -c_1 \sin x+?\]Ok good and derivative of sine gives us cosine yes?\[\Large y'(x)\quad=\quad -c_1 \sin x+c_2 \cos x\]

zepdrix (zepdrix):

So we've found the `general solution` and it's derivative:\[\Large y(x)\quad=\quad c_1\cos x+c_2 \sin x\]\[\Large y'(x)\quad=\quad -c_1 \sin x+c_2 \cos x\]

zepdrix (zepdrix):

We want to use our initial data to come up with a system of equations:\[\Large y(\pi/3)\qquad\to\qquad c_1\cos(\pi/3)+c_2\sin(\pi/3)\quad=\quad 0\]\[\Large y'(\pi/3)\qquad\to\qquad -c_1\sin(\pi/3)+c_2\cos(\pi/3)=2\]